### v1

# Testing for sufficient follow-up in survival data with a cure fraction

Abstract
In order to estimate the proportion of ‘immune’ or ‘cured’ subjects who will never experience failure, a sufficiently long follow-up period is required. Several statistical tests have been proposed in the literature for assessing the assumption of sufficient follow-up, meaning that the study duration is longer than the support of the survival times for the uncured subjects. However, for practical purposes, the follow-up would be considered sufficiently long if the probability for the event to happen after the end of the study is very small. Based on this observation, we formulate a more relaxed notion of ‘practically’ sufficient follow-up characterized by the quantiles of the distribution and develop a novel nonparametric statistical test. The proposed method relies mainly on the assumption of a non-increasing density function in the tail of the distribution. The test is then based on a shape constrained density estimator such as the Grenander or the kernel smoothed Grenander estimator and a bootstrap procedure is used for computation of the critical values. The performance of the test is investigated through an extensive simulation study, and the method is illustrated on breast cancer data.
1 Introduction
Cure models for analysis of time-to-event data in the presence of subjects who will never experience the event of interest has recently attracted increasing attention from both methodological and application perspectives. Survival data with a cure fraction are nowadays frequently encountered in oncology since advances in cancer treatments have resulted in a larger proportion of patients recovering from their illnesses and not experiencing relapse or cancer-related death (Legrand and Bertrand,, 2019). Among other fields, cure models have also been applied to loan credit scoring for analyzing the default time of a loan applicant, while a majority of debtors does not default (Dirick et al.,, 2017; Zhang et al.,, 2019). Independent of the application of interest, we refer to the subjects who are immune to the event of interest as ‘cured’ and to the susceptible ones as ‘uncured’. For a comprehensive review on cure models, we refer the reader to Maller and Zhou, (1996), Amico and Van Keilegom, (2018) and Peng and Yu, (2021).
In the presence of censoring, it is not possible to distinguish the cured subjects from the censored uncured ones. The presence of a cure fraction is however indicated by a Kaplan-Meier estimator (KME) of the survival function (Kaplan and Meier,, 1958) that reaches a plateau at a level greater than zero. The level of such plateau can be considered as an estimate of the cure fraction, provided that the follow-up is sufficiently long to assume that the survival function would remain constant even after the end of the study. In practice, it is however unclear what is the minimum required follow-up to accurately estimate the cure fraction and a KME with a prolonged plateau, which contains many censored observations, is considered as an indication of sufficient follow-up. Nevertheless, such visual inspection might often be ambiguous and inadequate for assessing sufficient follow-up, leading to an overestimation of the cure fraction. The crucial nature of this assumption underscores the necessity for a reliable statistical test.
The notion of ‘sufficient follow-up’ was first characterized in Maller and Zhou, (1992, 1994) as the setting in which the support of the event times for the uncured is included in the support of the censoring times. A procedure for testing the null hypothesis of insufficient follow-up, based on the length between the maximum observed time and the maximum uncensored event time, was introduced. Maller and Zhou, (1994) noticed that the Type I error probability of such procedure can be higher than the significance level and emphasized the need for theoretical analysis of the test statistic. In order to control the level of the test, an alternative approach was introduced by Maller and Zhou, (1996, p. 81). This method relies on simulations to approximate the distribution of the test statistics, by assuming the survival time of the susceptible and the censoring time have an exponential and a uniform distributions, respectively. To circumvent making parametric assumptions on the distributions, while still ensuring control over the level, Shen, (2000) introduced an alternative method based on the ratio of the two maximal event times, rather than their difference but the practical behavior of the test is still unsatisfactory. Klebanov and Yakovlev, (2007) pointed out that the test procedure by Maller and Zhou, (1994) is inadequate for testing sufficient follow-up also because of its non-monotonic behavior when the follow-up is lengthened. Recently, Maller et al., (2022) studied the finite sample and the asymptotic distributions of the two maximal censored and uncensored times. Such results were then used in Maller et al., (2023) to establish the asymptotic distribution of the test statistic in Maller and Zhou, (1996). On the other hand, Xie et al., (2023) considered testing the null hypothesis of sufficient follow-up and proposed a new test statistics that compares two nonparametric estimators of the cure rate, one from the KME and another based on an extrapolation of the KME using extreme value theory.
The existing characterization of sufficient follow-up requires that there exists a cutoff or ‘cure time’, such that surviving beyond such time is equivalent of being cured, and it is not possible for an event to happen after the end of the study. This is however not realistic and in practice follow-up would be considered sufficient if the chance of the event happening after the end of the study is negligible. For example, depending on the cancer type, relapse after 10 or 20 years is very rare but not impossible. In such cases, considering patients that survive 10 or 20 years relapse-free as cured and using the height of the plateau of the KME as estimator of the cure proportion would be adequate. Therefore we propose a more relaxed notion of sufficient follow-up which means that the probability for the event to happen after the end of the study is smaller than a prespecified threshold, e.g. 1%, chosen by the user. This would be more realistic in practice and still guarantee good identification of the cure proportion. Under the reasonable assumption that the survival time of the susceptible subjects has a non-increasing density in the tail region, one intuitively expects a small value of the density function at the end point of the support of the observed survival times, given the follow-up is sufficient. Therefore we utilize the Grenander or the kernel smoothed Grenander estimator of the density as our test statistic. The asymptotic properties of the test are studied, and its finite sample performance is investigated through an extensive simulation study. In terms of the level control and test power, the simulation study shows that the proposed method performs better compared to the existing tests for the null hypothesis of insufficient follow-up.
The article is organized as follows. Section 2 contains a description of existing formulations of testing sufficient follow-up along with a discussion on their differences. In Section 3, we propose a novel notion of ‘practically’ sufficient follow-up and introduce a procedure to test the null hypothesis of insufficient follow-up under the newly proposed characterization. The finite sample performance of the proposed method are investigated through a comprehensive simulation study in Section 4. To illustrate the use of the method in practice, we analyze two breast cancer datasets in Section 5. Proofs are presented in the Appendix, while additional simulation results can be found in the Supplementary Material. Software in the form of R code is available on the GitHub repository https://github.com/tp-yuen/cureSFUTest.
2 Problem formulation and discussion
Suppose $T$ is a nonnegative random variable denoting the event time of a subject with distribution function $F$ and density $f$. Let $C$ be the random right censoring time with distribution function $G$ and assume that $C$ and $T$ are independent. Under the random right censoring assumption, we observe the pair $\left(Y,\Delta\right)$ where $Y=\min{\left(T,C\right)}$ is the observed survival time and $\Delta=\mathds{1}_{\{{T\leq C}\}}$ is the censoring indicator. Let $H(t)=\mathbb{P}\left[Y\leq t\right]$ be the distribution of the observed survival time. In the presence of cured subjects in the population, the mixture cure model (MCM) assumes that a subject is susceptible with probability $p$ and is cured with probability $1-p$. With the convention that $\infty\cdot 0=0$, the survival time $T$ can be decomposed as $T=(1-B)\cdot\infty+B\cdot T_{u}$, where $B$ is a latent binary variable indicating the uncure status of a subject ($B=1$ if the subject is uncured) and $T_{u}$ is the survival time for the uncured subject with a proper distribution function $F_{u}(t)=\mathbb{P}\left[T\leq t|B=1\right]$. The subscript ‘u’ will be used to denote quantities that correspond to the uncured subpopulation. Then the distribution function of $T$ under the MCM is given by
$F(t)={\mathbb{P}\left[T\leq t,B=1\right]=\mathbb{P}\left[B=1\right]\mathbb{P}\left[T\leq t\mid B=1\right]}=pF_{u}(t),$
which is an improper distribution when $p<1$. Throughout this paper, we assume that $10<p<$ 1, indicating the presence of cured subjects in the population. The testing for $p=1$ against $p<1$ has been explored extensively in the literature, see Chapter 2 in Maller and Zhou, (1996), Section 2.1 in Shen, (2000), and Chapter 6 in Peng and Yu, (2021).
Let $\tau_{G}=\sup\left\{t\geq 0:G(t)<1\right\}$ and $\tau_{F_{u}}=\sup\left\{t\geq 0:F_{u}(t)<1\right\}$ be the right extremes of $G$ and $F$ respectively. Suppose that $n$ i.i.d. realizations $(Y_{i},\Delta_{i}),i=1,\dots,n$ of $(Y,\Delta)$ are observed. Based on the data, we can estimate nonparametrically the distribution $F$ by the KME (Kaplan and Meier,, 1958) defined as
$\hat{F}_{n}(t)=1-\prod_{i:Y_{i}\leq t}\left(1-\frac{\#\{j:Y_{j}=Y_{i},\Delta_{j}=1\}}{\#\{j:Y_{j}\geq Y_{i}\}}\right).$
Maller and Zhou, (1992) showed that, if $\tau_{G}\geq\tau_{F_{u}}$, the uncure fraction $p$ can be estimated consistently by $\hat{p}_{n}=\hat{F}_{n}(Y_{(n)})$, where $Y_{(n)}$ is the largest observed survival time. Otherwise, $\hat{p}_{n}$ would underestimate $p$. This gives rise to the crucial assumption of sufficient follow-up, i.e. that $\tau_{G}\geq\tau_{F_{u}}$. In practice $\tau_{F_{u}}$ is not known and a visual inspection of the KME is usually used to assess sufficient follow-up since under such assumption one expects to see a Kaplan–Meier estimate exhibiting a long plateau, which contains many censored observations. This paper aims to introduce a new statistical test for the sufficient follow-up assumption. First we discuss more extensively on the existing approaches mentioned in the intoduction section.
2.1 Testing insufficient versus sufficient follow-up
Maller and Zhou, (1992, 1994) were the first to propose a nonparametric test for testing
$H_{0}:\tau_{G}\leq\tau_{F_{u}}\quad\text{versus}\quad H_{a}:\tau_{G}>\tau_{F_{u}}.$ (1)
relying on the magnitude of $\tau_{G}-\tau_{F_{u}}$. Let $Y_{(n)}$ be the largest observed survival time and $\tilde{Y}_{(n)}$ be the largest event time (uncensored survival time). They showed that $Y_{(n)}-\tilde{Y}_{(n)}$ converges almost surely to $\tau_{G}-\tau_{F_{u}}$ if $\tau_{F_{u}}\leq\tau_{G}$ and to 00 if $\tau_{F_{u}}>\tau_{G}$, and suggested rejecting the null hypothesis $H_{0}$ if the $p$-value $\mathbb{P}[{Y_{(n)}-\tilde{Y}_{(n)}\geq y_{(n)}-\tilde{y}_{(n)}}]<\alpha$, where $\alpha$ is a pre-specified significance level, $y_{(n)}$ and $\tilde{y}_{(n)}$ are the observed values of $Y_{(n)}$ and $\tilde{Y}_{(n)}$, respectively. Under the null hypothesis $H_{0}$ and with a large sample size, the previous probability can be approximated reasonably by $(1-q_{n})^{n}$, where $q_{n}=\mathbb{P}[{C\geq{T>2\tilde{y}_{(n)}-{y}_{(n)}}}]$. Without knowing the distribution functions $F$ and $G$, one can estimate $q_{n}$ using its empirical estimator $\hat{q}_{n}$, which is given by
$\hat{q}_{n}=\frac{\sum_{i=1}^{n}\Delta_{i}\mathds{1}_{\{{2\tilde{y}_{(n)}-y_{(n)}<Y_{i}\leq\tilde{y}_{(n)}}\}}}{n}.$ (2)
Then $H_{0}$ is rejected if the estimated $p$-value $\alpha_{n}=\left(1-\hat{q}_{n}\right)^{n}$ is smaller than the nominal level. Following the convention of Maller and Zhou, (1996, p. 95), we call this test procedure the $\alpha_{n}$-test. Maller and Zhou, (1994, Section 5.3) noted that the Type I error probability of the $\alpha_{n}$-test can be larger than the significance level and advised resorting to simulations in such situation.
In order to control the Type I error for testing the hypotheses in (1), Shen, (2000) introduced an alternative method based on the $\alpha_{n}$-test, known as the $\tilde{\alpha}_{n}$-test. Rather than considering the difference between $Y_{(n)}$ and $\tilde{Y}_{(n)}$, the $\tilde{\alpha}_{n}$-test relies on the ratio of the two order statistics, and $H_{0}$ is rejected if the $p$-value $\mathbb{P}[{Y_{(n)}/\tilde{Y}_{(n)}>r_{n}}]<\alpha$, where $r_{n}=y_{(n)}/\tilde{y}_{(n)}$. The last probability is bounded by $\mathbb{P}[{\tilde{Y}_{(n)}<\tau_{G}/r_{n}}]$. Under $H_{0}$ and with a large $n$, this bound can be approximated by $\mathbb{P}[{\tilde{Y}_{(n)}<\hat{\tau}_{G}/r_{n}}]$, where $\hat{\tau}_{G}=\hat{w}\tilde{y}_{(n)}+(1-\hat{w})y_{(n)}$ with $\hat{w}=(y_{(n)}-\tilde{y}_{(n)})/y_{(n)}$. Similar to the $\alpha_{n}$-test, the empirical estimator ${\tilde{\alpha}}_{n}$ can be used, which is defined by
${\tilde{\alpha}}_{n}=\left(1-\frac{\sum_{i=1}^{n}\Delta_{i}\mathds{1}_{\{{\hat{\tau}_{G}/r_{n}\leq Y_{i}\leq\tilde{y}_{(n)}}\}}}{n}\right)^{n}.$ (3)
Shen, (2000) also investigated the Type I error and power of the $\tilde{\alpha}_{n}$-test and showed through simulations that despite not being optimal, the test controls the level better than the $\alpha_{n}$-test.
Maller et al., (2022) studied both finite sample and asymptotic distributions for the largest observed survival time $Y_{(n)}$ and the largest event time $\tilde{Y}_{(n)}$, and emphasized that such results can be used to determine the critical values of the previous tests. Based on the established results, Maller et al., (2023) further derived the finite sample and asymptotic distributions of the test statistics $n\hat{q}_{n}$, where $\hat{q}_{n}$ is defined in (2). When computing the finite-sample distribution of $n\hat{q}_{n}$, the distribution functions $F$ and $G$ play a crucial role and need to be estimated from the sample in practice. On the other hand, the asymptotic distribution of $n\hat{q}_{n}$ when the follow-up is sufficient/insufficient is obtained under some tail behavior assumptions on $F_{u}$ and $G$. These distributions can be used to determine the critical value for testing $H_{0}$ versus $H_{a}$ in (1). We refer to such test procedure as the $Q_{n}$-test. However, the simulation studies (Xie et al.,, 2023, Table 4) show that the empirical level of the $Q_{n}$-test is still larger than the nominal level and deviates more when the follow-up time increases.
2.2 Testing sufficient versus insufficient follow-up
Xie et al., (2023) considered testing the null hypothesis of sufficient follow-up
$\check{H}_{0}:\tau_{F_{u}}\leq\tau_{G}\quad\text{versus}\quad\check{H}_{a}:\tau_{F_{u}}>\tau_{G}.$ (4)
The authors proposed a test statistic $T_{n}$ given by the difference between an estimator of $p$ computed as if the follow-up was sufficient $\hat{F}_{n}(y_{(n)})$ and an estimator computed via extrapolation of $\hat{F}_{n}$ beyond $y_{(n)}$ based on extreme value theory. The extrapolation corrects the underestimation of $\hat{F}_{n}(y_{(n)})$ for $p$ under the insufficient follow-up setting. The asymptotic normality of the test statistic $T_{n}$ was established and a bootstrap procedure was introduced to approximate the critical values of the test, resulting in better approximation compared to the asymptotic results.
2.3 Flipping the hypotheses or not?
The difference between the two formulations of the hypotheses discussed previously is the type of error that one aims to control in practice. For the $\alpha_{n}$, $\tilde{\alpha}_{n}$ and $Q_{n}$ tests one aims to control the probability of concluding that the follow-up is sufficient when it is actually not, while $T_{n}$ controls the probability of deciding that the follow-up is insufficient when it is actually sufficient. From the practical point of view, if based on the plateau of the KME and medical knowledge, one expects that the follow-up is sufficient, the $T_{n}$ test for the null hypothesis of sufficient follow-up can be used. In this case sufficient follow-up would be rejected if there is evidence against it and otherwise one can proceed with estimation assuming sufficient follow-up. On the other hand, if one is uncertain about the sufficient follow-up assumption based on KME and medical knowledge, it is safer to test the null hypothesis of insufficient follow-up and conclude sufficient follow-up only if there is strong evidence in favor of it. This would be a safer choice since the consequences of assuming sufficient follow-up when it is actually not true are more serious than vice versa because it would lead to wrong conclusions regarding the cure fraction.
3 Testing procedure
In this paper we focus on testing the null hypothesis of insufficient follow-up. However our goal is to consider a more relaxed formulation of sufficient follow-up compared to the one in equation (1), which would be more realistic and satisfactory in practice. Note that in practice one always has $\tau_{G}<\infty$ because of the finite length of studies and as a result the characterization in equation (1) would mean that the follow-up is sufficient only if there exists a finite cure time $\tau_{F_{u}}$ such that it is impossible for an event to happen after $\tau_{F_{u}}$ and the study is longer than that. This is hardly ever the case in practical applications. For example, it is known that cancer relapse after 5, 10 or 20 years (depending on the cancer type) is very rare but yet not impossible and clinical trials are usually not very long to have the certainty that no events will happen after the end of the follow-up period. Hence, in practice it would be more meaningful to consider follow-up as sufficient when $F_{u}(\tau_{G})>1-\epsilon$ for some small $\epsilon>0$ (e.g. $\epsilon=0.01$), or equivalently, when $\tau_{G}>q_{1-\epsilon}$, with $q_{1-\epsilon}=\inf\left\{t\geq 0:F_{u}(t)\geq 1-\epsilon\right\}$. Specifically, we introduce the following hypotheses:
$\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}\quad\text{versus}\quad\tilde{H}_{a}:q_{1-\epsilon}<\tau_{G}.$ (5)
Note that under $\tilde{H}_{a}$, we have $F(\tau_{G})=pF_{u}(\tau_{G})\in[p-\epsilon p,p]$. Hence, the amount of underestimation of $p$, resulting from using $\hat{F}_{n}(\tau_{G})$ as an estimator, is very small (less than $\epsilon$). Such underestimation of $p$ is small relative to the estimation uncertainty for finite sample sizes and therefore negligible for practical purposes.
3.1 Idea of the test
Our test relies on the assumption that the density function $f_{u}$ of $F_{u}$ is non-increasing in the tail region and at least continuous. Then the hypothesis in (5) can equivalently be written as
$\tilde{H}_{0}:f(\tau_{G})\geq f(q_{1-\epsilon})\quad\text{versus}\quad\tilde{H}_{a}:f(\tau_{G})<f(q_{1-\epsilon}),$ (6)
since $f(t)=pf_{u}(t)$. Then the idea is to estimate $f(\tau_{G})$ from the sample with standard methods and to reject $\tilde{H}_{0}$ if such estimator is larger than a critical value. Note however that $f(q_{1-\epsilon})$ is unknown and we instead try to find a lower bound for that, which would guarantee the level of the test in the worse case scenario. The reasoning is as follows. Let $\eta>0$ be a very small number compared to $\epsilon$ and $\tau>q_{1-\epsilon}$ be such that $\mathbb{P}\left(T_{u}>\tau\right)<\eta$ or essentially $\mathbb{P}\left(T_{u}>\tau\right)$ is negligible. By the monotonicity and the continuity of $f_{u}$, we have
$\epsilon=\int_{q_{1-\epsilon}}^{\tau_{F_{u}}}f_{u}(t)\mathrm{d}{t}\leq\int_{q_{1-\epsilon}}^{\tau}f_{u}(t)\mathrm{d}{t}+\eta\leq f_{u}(q_{1-\epsilon})(\tau-q_{1-\epsilon})+\eta,$
meaning that under $\tilde{H}_{0}$
$f(\tau_{G})\geq f(q_{1-\epsilon})\geq\frac{(\epsilon-\eta)p}{\tau-\tau_{G}}\approx\frac{\epsilon p}{\tau-\tau_{G}}.$
Since $p\geq F(\tau_{G})$, the idea is to reject $\tilde{H}_{0}$ if an estimator $\hat{f}_{n}(\tau_{G})$ is smaller than $\epsilon\hat{F}_{n}(\tau_{G})/(\tau-\tau_{G})+\delta_{n}$, where $\delta_{n}$ is determined based on the desired level and the distribution of the estimator $\hat{f}_{n}(\tau_{G})$. More details are given below for two specific density estimators. The choice of $\tau$ can be based on some prior knowledge, for example, one can take $\tau>\tau_{G}$ such that it is almost impossible for the event to happen after $\tau$, i.e. if one had a follow-up of length $\tau$ one would consider it sufficient. As we illustrate in the simulation study and real data application below, one can also do a sensitivity analysis with respect to $\tau$ and, if in doubt, we suggest taking a larger $\tau$ for a more conservative test. Below we describe two possible approaches of estimation of $f(\tau_{G})$, leading to two different statistical tests. For simplicity we consider a non-increasing density on the whole support but the results hold more in general. In practice one would just need to restrict the shape constrained estimation on a subset of the support determined based on visual inspection of the KME (F is concave where $f$ is non-increasing) or a statistical test for monotonicity.
3.2 Grenander estimator of $f$
A natural nonparametric estimator of $f$ under monotonicity constraints would be the Grenander estimator (Grenander,, 1956) denoted by $\hat{f}_{n}^{G}$, which is defined as the left derivative of the least concave majorant (LCM) of the Kaplan-Meier estimator. We refer the reader to Groeneboom and Jongbloed, (2014) for a comprehensive overview of statistical inference techniques under shape-constraints. We are interested in estimating the density function at the boundary of the support of the observed data $\tau_{G}$ but it is well-known that $\hat{f}_{n}^{G}$ is not consistent at the boundaries (Woodroofe and Sun,, 1993). For the standard setting without censoring, Kulikov and Lopuhaä, (2006) investigated the behavior of $\hat{f}_{n}^{G}$ near the boundary of the support of $f$ and proposed using $\hat{f}_{n}^{G}(\tau_{G}-cn^{-a})$ as a consistent estimator of $f(\tau_{G})$. Here $c>0$ is a constant, and $10<a<$ 1 controls the rate of convergence and the limiting distribution of the estimator.
In the presence of random right censoring one can show, as in Kulikov and Lopuhaä, (2006), that the following holds. Here $W(t)$ denotes a Brownian motion and $D_{R}[Z(t)](a)$ is the right derivative of the LCM on $[0,\infty)$ of the process $Z(t)$ at the point $t=a$.
Theorem 1.
Assume that $f_{u}$ is nonincreasing and differentiable with bounded derivative on $[0,\tau_{G}]$, $\tau_{G}<\infty$. If $f(\tau_{G})>0$, $G(\tau_{G}-)<1$ and $c>0$ is a fixed constant, we have
  • i)
    for $a\in(1/3,1)$ and $A_{1}=\sqrt{c[1-G(\tau_{G}-)]/f(\tau_{G})}$, the sequence
    $A_{1}n^{(1-a)/2}\left\{f(\tau_{G})-\hat{f}_{n}^{G}(\tau_{G}-cn^{-a})\right\}$
    converges in distribution to $D_{R}[W(t)](1)$ as $n\to\infty$.
  • ii)
    the sequence
    $A_{2}n^{1/3}\left\{f(\tau_{G})-\hat{f}_{n}^{G}(\tau_{G}-cB_{2}n^{-1/3})\right\}$
    converges in distribution to $D_{R}[W(t)-t^{2}](c)$ as $n\to\infty$, where
    ${B_{2}}=\left(\frac{2\sqrt{f(\tau_{G})}}{|f^{\prime}(\tau_{G})|\sqrt{1-G(\tau_{G}-)}}\right)^{2/3}\quad\text{and}\quad{A_{2}}=\sqrt{\frac{B_{2}[1-G(\tau_{G}-)]}{f(\tau_{G})}}.$
Note that the assumption $G(\tau_{G}-)<1$ indicates that the censoring distribution has a positive mass at $\tau_{G}$ and is required in order to have a strong approximation of the KME by a Brownian motion appearing in the limiting distribution. As this mass becomes smaller, the constants $A_{2}$ and $B_{2}$ get closer to zero and infinity respectively. Such assumption is reasonable since one can think of the censoring variable as $C=\min(\tilde{C},\tau_{G})$, where $\tau_{G}$ represents the study duration (administrative censoring), while $\tilde{C}$ is the censoring time due to loss of follow-up for other reasons and would take even values larger than $\tau_{G}$ had the study continued for longer. The previous theorem shows that the $\hat{f}_{n}^{G}(\tau_{G}-cB_{2}n^{-1/3})$ converges at rate $n^{1/3}$ to $f(\tau_{G})$ but computation of $B_{2}$ requires also estimation of the derivative of the density at $\tau_{G}$. To avoid that, for the test statistics, we will instead use an estimator $\hat{f}_{n}^{G}(\tau_{G}-cn^{-a})$ for some $a$ larger than but close to $1/3$ as in statement (i) of the theorem. Such choice, despite being sub-optimal in terms of rate of convergence, behaves better in practice.
3.3 Smoothed Grenander estimator
Alternatively, we can consider the kernel smoothed Grenander-type estimator of $f(\tau_{G})$. Let $k$ be a standard kernel function, which is a symmetric probability density function with support $[-1,1]$, and $h$ be the bandwidth parameter. The triweight kernel is used throughout this work, i.e. $k(v)=\frac{35}{32}(1-v^{2})^{3}\mathds{1}_{[-1,1]}(v)$. To circumvent the inconsistency issues of the standard kernel density estimator at the boundary, various boundary correction methods have been proposed. One such approach utilizes boundary kernels, which replaces the standard kernel near the boundary with linear combinations of $k(v)$ and $vk(v)$, with coefficients depending on the value of $k$ near the boundary (Zhang and Karunamuni,, 1998; Durot et al.,, 2013). Denoting this boundary kernel as $k_{B,t}(v)$, which is defined by
$k_{B,t}(v)=\begin{cases}\phi\left(\frac{t}{h}\right)k(v)+\psi\left(\frac{t}{h}\right)vk(v)&\quad t\in[0,h],\\ k(v)&\quad t\in(h,\tau_{G}-h),\\ \phi\left(\frac{t}{h}\right)k(v)-\psi\left(\frac{t}{h}\right)vk(v)&\quad t\in[\tau_{G}-h,\tau_{G}].\end{cases}$
Here, for $s\in[-1,1]$, the coefficients $\phi(s)$ and $\psi(s)$ are determined by
$\begin{split}\phi(s)\int_{-1}^{s}k(v)\mathrm{d}{v}&+\psi(s)\int_{-1}^{s}vk(v)\mathrm{d}{v}=1,\\ \phi(s)\int_{-1}^{s}vk(v)\mathrm{d}{v}&+\psi(s)\int_{-1}^{s}v^{2}k(v)\mathrm{d}{v}=0.\end{split}$
The smoothed Grenander-type estimator with boundary correction of $f(\tau_{G})$ is given by:
$\hat{f}_{nh}^{SG}(\tau_{G})=\int_{\tau_{G}-h}^{\tau_{G}}\frac{1}{h}k_{B,\tau_{G}}\left(\frac{\tau_{G}-v}{h}\right)\hat{f}_{n}^{G}(v)\mathrm{d}{v}.$ (7)
Such estimator and its asymptotic distribution in the interior of the support has been studied in Lopuhaä and Musta, (2017). Here we extend such results to the estimation at the right boundary $\tau_{G}$. In the following, let $k$ be a symmetric twice continuously differentiable kernel with support $[-1,1]$ such that $\int k(u)\mathrm{d}{u}=1$ and its first derivative is bounded uniformly.
Theorem 2.
Assume that $f_{u}$ is a nonincreasing and twice continuously differentiable density with $f_{u}$ and $|f_{u}^{\prime}|$ being strictly positive. If $G(\tau_{G}-)<1$ and $hn^{1/5}\xrightarrow{}c\in(0,\infty)$ as $n\xrightarrow{}\infty$, then
[…]
where
$\mu=\frac{1}{2}c^{2}f^{\prime\prime}(\tau_{G})\int_{0}^{1}v^{2}k_{B,\tau_{G}}(v)\mathrm{d}{v},\quad\text{and}\quad\sigma^{2}=\frac{f(\tau_{G})}{c\left[1-G(\tau_{G}-)\right]}\int_{0}^{1}k_{B,\tau_{G}}(v)^{2}\mathrm{d}{v}.$
Note that the rate of convergence of the smooth Grenander estimator is faster than the non-smoothed one and the limiting variance becomes larger as the mass of the censoring distribution at $\tau_{G}$ becomes smaller. The limiting bias depends on the second derivative of the density which is difficult to be estimated so in practice we will use a bootstrap procedure to approximate the critical value of the test as described later.
3.4 Test procedure based on asymptotic results
Based on the idea described in Section 3.1 and the two estimators of a decreasing density, we propose two statistical tests for the assumption of ‘practically’ sufficient follow-up formulated in (5). The first test rejects the null hypothesis of insufficient follow-up $\tilde{H}_{0}$ if
$\hat{f}_{n}^{G}(\tau_{G}-cn^{-a})\leq\frac{\epsilon\hat{F}_{n}(\tau_{G})}{\tau-\tau_{G}}-A_{1}^{-1}n^{-(1-a)/2}Q_{1-\alpha}^{G},$ (9)
where $a\in(1/3,1)$ (we choose $a=0.34$ in the simulation study, which is very close to 1/3), $c=\tau_{G}$ (so that the amount of deviation from the right boundary scales with the length of the support), $A_{1}$ is defined in Theorem 1 and $Q_{\alpha}^{G}$ is the $\alpha$-quantile of the distribution $D_{R}[W(t)](1)$. Below we show that the level of the test is asymptotically bounded by $\alpha$.
Proposition 1.
Suppose $0<\eta<\epsilon$ is such that $n^{(1-a)/2}\eta\to 0$ as $n\to\infty$. Then, for any $\tau_{G}\leq q_{1-\epsilon}$,
$\lim_{n\to\infty}\mathbb{P}\left[\hat{f}_{n}^{G}(\tau_{G}-cn^{-a})\leq\frac{\epsilon\hat{F}_{n}(\tau_{G})}{\tau-\tau_{G}}-A_{1}^{-1}n^{-(1-a)/2}Q_{1-\alpha}^{G}\right]\leq\alpha.$
The advantage of this test is that it does not require smoothing and twice differentiability of the true density function. However, it requires a plug-in estimate for $A_{1}$ and a choice of the constant $c$. In practice we observe that very large sample sizes would be required to have a good performance of this test.
The second test we propose rejects the null hypothesis of insufficient follow-up $\tilde{H}_{0}$ if
$\hat{f}_{nh}^{SG}(\tau_{G})\leq\frac{\epsilon\hat{F}_{n}(\tau_{G})}{\tau-\tau_{G}}+n^{-2/5}Q_{\alpha}^{SG},$ (10)
where $Q_{\alpha}^{SG}$ denotes the $\alpha$-quantile of the distribution $N(\mu,\sigma^{2})$ in Theorem 2. Such test is also guaranteed to be asymptotically correct.
Proposition 2.
Suppose $0<\eta<\epsilon$ is such that $n^{2/5}\eta\to 0$ as $n\to\infty$. Then, for any $\tau_{G}\leq q_{1-\epsilon}$,
$\lim_{n\to\infty}\mathbb{P}\left[\hat{f}_{nh}^{SG}(\tau_{G})\leq\frac{\epsilon\hat{F}_{n}(\tau_{G})}{\tau-\tau_{G}}+n^{-2/5}Q_{\alpha}^{SG}\right]\leq\alpha.$
Computation of the quantiles $Q_{\alpha}^{SG}$ requires plug-in estimates for the asymptotic mean and variance of the smooth Grenander estimator, which do not behave very well for small sample sizes. For this reason, we suggest a bootstrap procedure for approximating the critical value of the test as describe below.
3.5 Bootstrap procedure
It is worthy to mention that the naive bootstrap, by resampling with replacement from the pairs $(y_{i},\delta_{i})$ and computing the Grenander estimator $\hat{f}_{n}^{G}(\tau_{G})$ from the bootstrap samples, is inconsistent, since the true density of the bootstrap sample is not continuous (Kosorok,, 2008; Sen et al.,, 2010; Sen and Xu,, 2015). Therefore a smoothed bootstrap procedure, based on the smoothed Grenander estimator $\hat{f}_{n}^{SG}$, is introduced to approximate the critical value of the test and is delineated in Algorithm S1 in the Supplementary Material. The idea of the smoothed bootstrap procedure is that a modified kernel density estimator $\tilde{f}_{nh_{0}}$ is estimated from the sample data, followed by drawing bootstrap samples from $\tilde{f}_{nh_{0}}$. Specifically, $\tilde{f}_{nh_{0}}$ is the derivative of the kernel smoothed least concave majorant $\hat{F}^{G}_{n}$ of the KME, i.e. the derivative of $\int\frac{1}{h_{0}}k_{B,t}({\frac{t-u}{h_{0}}})\hat{F}_{n}^{G}(u)\mathrm{d}{u}$. If one would use the bandwidth $h$ of order $n^{-2/5}$ as in the computation of $\hat{f}_{nh}^{SG}$, an explicit estimation of $f^{\prime\prime}(\tau_{G})$ would be required to adjust the asymptotic bias of the bootstrap estimates (Groeneboom and Hendrickx,, 2017, 2018). Hence obtaining an accurate estimate for $f^{\prime\prime}(\tau_{G})$ would be crucial, see for example Section 2.5 in Groeneboom and Hendrickx, (2018), but it is problematic especially near the end point $\tau_{G}$. Instead, as proposed by Groeneboom and Jongbloed, (2023) in the monotone regression setting, we draw bootstrap samples from $\tilde{f}_{nh_{0}}$ with $h_{0}=0.7\tau_{G}n^{-1/9}\wedge 0.5\tau_{G}$, which leads to an over-smoothed estimate. Oversmoothing circumvents the need to correct for the asymptotic bias. Then we replace $Q^{SG}_{\alpha}$ in the critical value of the test in (10) by the $\alpha$-quantile of the bootstrap estimates $n^{2/5}\{\hat{f}_{nh}^{SG^{\ast}}(\tau_{G})-\tilde{f}_{nh_{0}}(\tau_{G})\}$.
4 Simulation study
In this section, we study the finite sample performance of the testing procedure (using $\hat{f}_{n}^{G}$ and $\hat{f}_{nh}^{SG}$) described in Section 3 for testing $\tilde{H}_{0}:q_{1-\epsilon}\leq\tau_{G}$ and compare it with the $\alpha_{n}$, $\tilde{\alpha}_{n}$ and $Q_{n}$ tests. In order to cover various scenarios with different uncured survival time and censoring time distributions, we consider six settings as described below. In Settings 13 the survival time of the uncured has unbounded support, i.e. $\tau_{F_{u}}=\infty$, while that of the censoring time is bounded, i.e. $\tau_{G}<\infty$. Such a scenario is the main focus of our methodology but it means that the follow-up is never sufficient based on the characterization of sufficient follow-up in (1) used by the existing tests. Settings 45 correspond to having $\tau_{F_{u}}<\infty$ and $\tau_{G}<\infty$, indicating that the follow-up is sufficient, under the notion in (1), when $\tau_{F_{u}}<\tau_{G}$. Setting 6 corresponds to having $\tau_{F_{u}}=\infty$ and $\tau_{G}=\infty$. Such case was considered by Xie et al., (2023) and is regarded as sufficient follow-up under the characterization in (4). The $T_{n}$ test of Xie et al., (2023) considers a null hypothesis of sufficient follow-up and as a result cannot be compared with ours. However, for this last setting we also present the results of the $T_{n}$ test.
4.1 Simulation settings
For Settings 13 and 6, three different uncured fractions $p$ are considered, namely 0.2, 0.6 and 0.8, to study the effect of $p$ on the testing performance. Note that the censoring rate is at least $1-p$ since all the cured subjects are observed as censored. For settings 4 and 5, the uncured rate $p$ is 0.6. We consider distribution that has a decreasing density for the uncured event time $T_{u}$. We generate the censoring time $C$ as $C=\min(\tilde{C},\tau_{G})$, where the support of $\tilde{C}$ includes $[0,\tau_{G}]$. In Section 3, the asymptotic properties of both the Grenander $\hat{f}_{n}^{G}$ and the smooth Grenander $\hat{f}_{nh}^{SG}$ estimators are affected by the mass of the censoring distribution at $\tau_{G}$, denoted by $\Delta G(\tau_{G})=1-G(\tau_{G}-)$. To investigate the effect of $\Delta G(\tau_{G})$ on the performance of the test, we consider $\tilde{C}$ with a uniform distribution on $[0,\zeta]$ for $\zeta\geq\tau_{G}$ chosen in such a way that $\Delta G(\tau_{G})\in\left\{0,0.02,0.05,0.2\right\}$, in settings 1, 3 and 4. For such settings, we have that the higher the mass $\Delta G(\tau_{G})$, the lower the censoring rate. For settings 15, 12 different $\tau_{G}$’s are considered in order to investigate the level and the power of the test. We choose the values of the $\tau_{G}$ in terms of quantiles of $F_{u}$ because, in practice, the probability of the event to happen after the end of the study is a more important measure of deviation from sufficient follow-up compared to the value of $\tau_{G}$. Specifically, we choose $q_{1},q_{2},q_{3},q_{4},q_{6},q_{12}$ to be the 90%, 92.5%, 95%, 97.5%, 99% and 99.9% quantiles of $F_{u}$. In addition we consider $q_{5}$ to be the mid-point between the 97.5% and 99% quantiles of $F_{u}$ and $q_{7}-q_{11}$ to be five evenly separated points between the 99% and 99.9% quantiles of $F_{u}$. Table S6 in the Supplementary Material reports their actual values for settings 15. If we choose $\epsilon=0.01$ for testing $\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}$ versus $\tilde{H}_{a}:q_{1-\epsilon}<\tau_{G}$, the cases when $q_{0.99}\geq\tau_{G}$ are considered as insufficient follow-up. Conversely, the follow-up for the cases when $q_{0.99}<\tau_{G}$ is sufficient. Therefore an ideal testing procedure is expected to give low rejection rate for the former case and high rejection rate for the latter case.
  1. Setting 1.
    The uncured subjects have the exponential distribution with rate parameter 1. The censoring times are generated using the aforementioned uniform distribution for $\tilde{C}$. For different choices of $\tau_{G}$, the censoring rate ranges from 48% to 64% (30%–52%) when $p=0.6$ (0.8) and $\Delta G(\tau_{G})=0.02$.
  2. Setting 2.
    The uncured have the exponential distribution with rate parameter $\lambda\in\{0.4,1,5\}$. The censoring times are generated using $\tilde{C}$ having exponential distribution with rate parameter 0.5. In this case $\Delta G(\tau_{G})=1-\exp(-0.5\tau_{G})$ and the higher the rate parameter $\lambda$, the lower the censoring rate. For different choices of $\tau_{G}$, the censoring rate ranges from 45% to 50% (60%–62%) when $p=0.6$ and $\lambda=5$ (1).
  3. Setting 3.
    The uncured subjects have the Weibull distribution with shape and scale parameters 0.5 and 1.5, respectively. Such distribution for the uncured has a density function that decreases faster comparing to the exponential distribution in Setting 1.The censoring times are generated using the aforementioned uniform distribution for $\tilde{C}$. For different choices of $\tau_{G}$, the censoring rate ranges from 42% to 56% (42%–53%) when $p=0.6$ and $\Delta G(\tau_{G})=0$ (0.2).
  4. Setting 4.
    The uncured have the truncated exponential distribution with rate parameter 1 and endpoint $\tau_{F_{u}}$, where $\tau_{F_{u}}$ is the 99% quantile of the exponential distribution with rate parameter 1, i.e. $\tau_{F_{u}}\approx 4.6$. The censoring times are generated using the aforementioned uniform distribution for $\tilde{C}$. To study the testing performance when $\tau_{F_{u}}\leq\tau_{G}<\infty$, for this setting we consider few extra $\tau_{G}$’s, which are greater than $\tau_{F_{u}}$, in addition to the $\tau_{G}$’s mentioned previously. For different choices of $\tau_{G}$, the censoring rate ranges from 48% to 64% (46%–61%) when $p=0.6$ and $\Delta G(\tau_{G})=0$ (0.2).
  5. Setting 5.
    The uncured have the truncated exponential distribution with rate parameter 5 and endpoint $\tau_{F_{u}}$, where $\tau_{F_{u}}$ is the 99% quantile of the exponential distribution with rate parameter 5, i.e. $\tau_{F_{u}}\approx 0.92$. The censoring times are generated using $\tilde{C}$ having the exponential distribution with rate parameter $\lambda_{C}\in\{0.5,3\}$. In this case $\Delta G(\tau_{G})=1-\exp(-\lambda_{C}\tau_{G})$ and the higher the rate parameter for the censoring distribution $\lambda_{C}$, the higher the censoring rate. As in Setting 4, we include extra $\tau_{G}$’s, which are greater than $\tau_{F_{u}}$, to investigate the testing performance when $\tau_{F_{u}}\leq\tau_{G}<\infty$. For different choices of $\tau_{G}$, the censoring rate ranges from 45% to 50% (62%–63%) when $p=0.6$ and $\lambda_{C}=0.5$ (3).
  6. Setting 6.
    The uncured have the exponential distribution with rate parameter 1. The censoring time follows the exponential distribution with rate parameter 0.5. The censoring rates, when $p=0.2$, 0.6 and 0.8, are around 86%, 60% and 46%, respectively.
The range of the censoring rates over different values of $\tau_{G}$ for all the considered settings are reported in Tables S1S5 in the Supplementary Material. The censoring rate depends on $p$, $\Delta G(\tau_{G})$ in settings 1, 3 and 4, $\lambda$ in setting 2, and $\lambda_{C}$ in setting 5 as mentioned previously.
For the proposed method, $\epsilon$ is set at 0.01, meaning that we consider the follow-up as sufficient when $q_{0.99}<\tau_{G}$. In Setting 3, we also investigate how the methods performs when $\epsilon=0.005$, resulting in a more conservative hypothesis of ‘practically’ sufficient follow-up. The parameter $\tau$ is set to the 99.95% quantile of the uncured survival time distribution $F_{u}$ for Settings 15. A different choice is made for Setting 6 since $\tau_{G}=\infty$. We also studied the effect of the choice of $\tau$ on the rejection rate, using Settings 1 and 4. The parameters $a$ and $c$ of the Grenander estimator $\hat{f}_{n}^{G}$ are set to 0.34 and $\tau_{G}$ ($y_{(n)}$ for Setting 6), respectively. For the smoothed Grenander estimator $\hat{f}_{nh}^{SG}$, the tri-weight kernel is used together with a bandwidth of $h=\tau_{G}n^{-1/5}\wedge 0.5\tau_{G}$ ($\tau_{G}$ is replaced by $y_{(n)}$ in Setting 6). We use 500 bootstrap samples to compute the critical value of the test based on the smoothed Grenander estimator as described in Section 3.5. The parameter $\gamma$ of the asymptotic distribution of the test statistic $n\hat{q}_{n}$ for the $Q_{n}$-test is set to 1, which covers a wide range of censoring distributions including those in the current simulation settings. We use the same setting as described in Section 4 of Xie et al., (2023) for the $T_{n}$-test with 500 bootstrap samples. The significance level of the tests is set at $0.05$.
4.2 Simulation results
For each simulation setting, three sample sizes are considered, namely 250, 500 and 1000, and for each of them we consider 500 replications. For Settings 15, we plot the rejection rate of insufficient follow-up against $\tau_{G}$ for different methods. For Setting 6, which corresponds to only one choice $\tau_{G}=\infty$, the rejection rate is reported in a table rather than a plot. In this section we only present the results for settings 1 and 4 and compare the two proposed tests to the $Q_{n}$ test, which is the best of the existing tests for the null hypothesis of insufficient follow-up. The rest of the results and comparison also with the other methods can be found in the Supplementary Material.
Figure 1: Rejection rate of the null hypothesis of insufficient follow-up for different methods (solid: $\hat{f}_{nh}^{SG}$, dashed: $\hat{f}_{n}^{G}$, dotted: $Q_{n}$) in Setting 1 with $n=500$, $\Delta G(\tau_{G})=0.02$ and $p$=0.2 (left), $p=0.6$ (center), $p=0.8$ (right).
Figure 1 depicts the rejection rate of the null hypothesis of insufficient follow-up against $\tau_{G}$ for Setting 1 when $n=500$, $\Delta G(\tau_{G})=0.02$ and for different choices of the uncured fraction $p$. In each subplot, $q_{1}$, …, $q_{12}$ refer to the 12 different $\tau_{G}$’s considered as described in the simulation settings. The horizontal dash-dotted line at 0.05 indicates the significance level $\alpha$, and the vertical one at $q_{6}$ indicates $\tau_{G}=q_{0.99}$ (the 99% quantile of $F_{u}$). Recall that we are testing the null hypothesis $\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}$ with $\epsilon=0.01$ and an ideal testing procedure is expected to reject the null less when $\tau_{G}<q_{0.99}$ (left-side of the vertical dash-dotted line) while to reject more when $\tau_{G}>q_{0.99}$ (right-side of the vertical dash-dotted line). The rejection rates of $\hat{f}_{nh}^{SG}$, $\hat{f}_{n}^{G}$, and $Q_{n}$ tests are represented by solid, dashed, and dotted lines, respectively, in each subplot. Each of these lines is constructed by interpolating the rejection rates between two consecutive $q$’s.
From Figure 1 we see that the proposed method using $\hat{f}_{nh}^{SG}$ performs better, in terms of level empirically, among the three procedures for different $p$’s, while the test using $\hat{f}_{n}^{G}$ is the worst. The poor performance of $\hat{f}_{n}^{G}$ in controlling the level probably results from the poor approximation of the critical value using the asymptotic distribution at a relatively small sample size. In terms of the empirical power, $\hat{f}_{nh}^{SG}$ behaves better than the $Q_{n}$ test in general, although there is a region for $\tau_{G}$ near $q_{6}$ where the $Q_{n}$ test has higher power. For the effect of $p$, we observe that as $p$ increases the performance improves, i.e. better control of the level and higher power. Note that when $p$ is small the censoring rate is very high (around 85% for $p=0.2$) and it becomes more difficult to detect that the follow-up is not sufficient. In such cases, larger sample sizes and $\Delta G(\tau_{G})$ are required in order to control the level of the test (see Figure S1 in the Supplementary Material). We also point out that for the formulation of sufficient follow-up in (1) considered by the $Q_{n}$ test, the null hypothesis is always true for this setting, but the observed rejection rate is far larger than the nominal level of $5\%$.
Figure 2: Rejection rate of the null hypothesis of insufficient follow-up for different methods (solid: $\hat{f}_{nh}^{SG}$, dashed: $\hat{f}_{n}^{G}$, dotted: $Q_{n}$) in Setting 1 with $n=500$, $p=0.6$ and $\Delta G(\tau_{G})$=0 (left), 0.02 (center), and 0.2 (right).
Figure 2 shows the rejection rate against $\tau_{G}$ when $n=500$ and $p=0.6$ are fixed, while changing $\Delta G(\tau_{G})$. The proposed procedure using $\hat{f}_{nh}^{SG}$ possesses a better control on the empirical level at each $\Delta G(\tau_{G})$, among the three methods, while the Type I error probability of $\hat{f}_{n}^{G}$ remains higher than the others empirically. The test based on $\hat{f}_{nh}^{SG}$ also has a higher rejection rate under the alternative hypothesis $\tilde{H}_{0}:\tau_{G}<q_{1-\epsilon}$, apart from in a small region near $q_{1-\epsilon}$. The performance of all tests improves, in terms of empirical level, as $\Delta G(\tau_{G})$ increases. We note that the asymptotic distributions of the estimators $\hat{f}_{n}^{G}$ and $\hat{f}_{nh}^{SG}$ rely on the assumption that $\Delta G(\tau_{G})>0$. The simulation results suggest the proposed method still works well when such assumption is violated, i.e. $\Delta G(\tau_{G})=0$.
Figure 3: Rejection rate of the null hypothesis of insufficient follow-up for different methods (solid: $\hat{f}_{nh}^{SG}$, dashed: $\hat{f}_{n}^{G}$, dotted: $Q_{n}$) in Setting 1 with $p=0.6$, $\Delta G(\tau_{G})=0.02$ and a sample size of 200 (left), 500 (center), and 1000 (right).
In Figure 3 we investigate the effect of the sample size, for fixed $p=0.6$ and $\Delta G(\tau_{G})=0.02$. As expected, we observe a better control of the level and a steeper power curve when the sample size is larger. The proposed procedure using $\hat{f}_{nh}^{SG}$ performs better than others for all three different sample sizes. Complete results for all combinations of $p$, $\Delta G(\tau_{G})$ and $n$ can be found in Figures S1-S3 in the Supplementary Material. In terms of sensitivity of the test with respect to the choice of $\tau$, we observe in Figure S4 that as $\tau$ increases, the rejection rate decreases, resulting in a more conservative test. The effect is more visible in terms of power than level.
Similar conclusions can be drawn from the simulation results in Settings 2-3 reported in the Supplementary Material. We point out that when there is strong censoring and the uncured fraction is small, all methods exhibited very poor performance in terms of empirical level. Problems occur also when the density $f_{u}$ decreases faster, as in Setting 3, for which the tests have a better control of the level when $p$ and $n$ are larger. For a more conservative test, one can consider a smaller $\epsilon$, which results in lower rejection rates as illustrated in Figure S11, but the difference is not very large.
Figure 4: Rejection rate of the null hypothesis of insufficient follow-up for different methods (solid: $\hat{f}_{nh}^{SG}$ with $\tau_{1}\approx 4.5569$, dashed: $\hat{f}_{nh}^{SG}$ with $\tau_{2}=1.25\tau_{1}$, dotted: $\hat{f}_{nh}^{SG}$ with $\tau_{2}=1.5\tau_{1}$, dash-dotted: $Q_{n}$) in Setting 4 with $n=500$, $p=0.6$ and $\Delta G(\tau_{G})$=0 (left), $\Delta G(\tau_{G})$=0.02 (center), and $\Delta G(\tau_{G})$=0.2 (right).
Next we consider Setting 4, in which $f_{u}$ has compact support. Figure 4 depicts the rejection rate of insufficient follow-up against $\tau_{G}$ when $n=500$ and $p=0.6$ are fixed, while varying $\Delta G(\tau_{G})$. In each subplot, $q_{1}$, …, $q_{12}$ refer to the 12 $\tau_{G}$’s mentioned in the simulation settings. We introduce 7 extra $\tau_{G}$’s ($q_{13}$ to $q_{19}$ ranging from $\tau_{F_{u}}\approx 4.6051$ to 6.9078) to investigate the power of the test. The vertical dash-dotted line indicates $\tau_{G}=q_{0.99}$ (the 99% quantile of $F_{u}$), and the horizontal one indicates the significance level $\alpha$. We considered 3 different $\tau$’s, namely $\tau_{1}$ being the 99.95% quantile of $F_{u}$ as described in the previous section, $\tau_{2}=1.25\tau_{1}$, and $\tau_{3}=1.5\tau_{1}$. The rejection rates of $\hat{f}_{nh}^{SG}$ using $\tau_{1}$, $\tau_{2}$, $\tau_{3}$, and the $Q_{n}$-test are rendered by solid, dashed, dotted, and dash-dotted lines, respectively, in each subplot. Each line is constructed by interpolating the rejection rate between two successive $q$’s. We again observe that as $\tau$ increases, the test becomes more conservative. For $\tau=\tau_{1}$, the rejection rate starts being higher than the nominal level of $5\%$ when $\tau_{G}>q_{4}=q_{0.975}$, which is still reasonable in practice, while the $Q_{n}$ test has higher rejection rate even for shorter follow-up. For $\tau_{2}$ and $\tau_{3}$ the test behaves well in terms of level and has most of the time has also higher power compared to $Q_{n}$. Overall, we advice to take $\tau$ possibly larger than $\tau_{F_{u}}$ when it is believed that the $F_{u}$ has compact support based on practical knowledge of an approximate value for $\tau_{F_{u}}$. Similar results are observed in Setting 5, see Figure S13. However, in that case the test $Q_{n}$ exhibits a good control of the level but has almost no power in detecting sufficient follow-up.
5 Real data application
In this section we illustrate the practical application of the proposed method and compare it with the existing tests ($\alpha_{n}$, $\tilde{\alpha}_{n}$, $Q_{n}$ and $T_{n}$) through two breast cancer datasets, one with a sample size of 278 and another with 1233. Besides using the data with a follow-up cutoff at the end of the study, we construct ‘what-if’ scenarios based on the same dataset in which the follow-up cutoff is earlier than the actual one. In particular, the subjects with the original follow-up time greater than the hypothetical cutoff are considered as censored with a follow-up time equal to the hypothetical cutoff time. Such scenarios help explore whether a testing procedure possesses a monotonicity behavior of deciding sufficient follow-up if the follow-up period is lengthened as considered in Klebanov and Yakovlev, (2007).
For the proposed procedure (using $\hat{f}_{n}^{G}$ and $\hat{f}_{nh}^{SG}$), the right extreme of the censoring distribution $\tau_{G}$ is approximated by the maximum observed survival time $y_{(n)}$. We used two choices of $\epsilon$, namely 0.01 and 0.025, indicating the follow-up is considered as sufficient when $\tau_{G}>q_{0.99}$ for the former case and similarly for the latter. The distribution of $\hat{f}_{nh}^{SG}$ is approximated using the bootstrapping procedure described in Algorithm S1 with 1000 bootstrap samples and bandwidths as mentioned in Section 4. We set the constants $a$ and $c$ to 0.34 and $y_{(n)}$, respectively, for the procedure using the Grenander estimator $\hat{f}_{n}^{G}$. The parameter $\gamma$ for the $Q_{n}$ is set at $1$ as in Section 4. The critical value of the $T_{n}$ test is computed using 1000 bootstrap samples.
5.1 Breast cancer study I
We analyze a dataset of an observational study consisting of 286 lymph-node-negative breast cancer patients. The patients received treatment between 1980 and 1995, whose age at treatment ranges form 26 to 83 years (with a median of 52 years). The relapse-free survival time (in months), i.e. the time until death or occurrence of distant metastases, is considered. The uncensored survival time ranges from 2 to 80 months, and the censoring rate is around 62.59%. The same dataset has been studied by Amico et al., (2019) and Dobler and Musta, (2023) under the mixture cure model settings. We refer the reader to Wang et al., (2005) for a more detailed description of the data.
Figure 5: Kaplan-Meier estimate (solid) and its least concave majorant (dotted) for the breast cancer observational study data.
Figure 5 shows the Kaplan-Meier curve, which exhibits a long plateau. The maximum uncensored survival time $\tilde{y}_{(n)}$ and the maximum observed survival time $y_{(n)}$ are 80 and 171 months, respectively, and there are 152 censored observations between $\tilde{y}_{(n)}$ and $y_{(n)}$ (i.e. on the plateau). A graphical inspection based on the plateau of the Kaplan-Meier curve suggests the follow-up is sufficient. In addition, the Kaplan-Meier curve supports the assumption of decreasing density in the tail (the curve is close to concave at least starting from around $25$ months). One could use some data-driven method to estimate the point at which the density starts being decreasing, but that is not crucial here since it will not affect the estimator at the end point. We apply the proposed testing procedure (using $\hat{f}_{n}^{G}$ and $\hat{f}_{nh}^{SG}$) and the other existing methods ($\alpha_{n}$, $\tilde{\alpha}_{n}$, $Q_{n}$ and $T_{n}$) for the whole follow-up and the hypothetical scenarios with shorter follow-up as mentioned at the beginning of this section. For the proposed testing procedure (using $\hat{f}_{n}^{G}$ and $\hat{f}_{nh}^{SG}$), we used two choices of $\tau$, namely 240 and 360 months, meaning that we are considering the probability of relapse after 20 or 30 years to be almost zero.
Table 1: $p$-values of testing $\tilde{H}_{0}$, $H_{0}$ or $\check{H}_{0}$ at different follow-up cutoffs for the breast cancer observational study (rounded to three decimal places). Bold indicates rejecting the null at the 5% level.
Cutoff $\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}$ $H_{0}:\tau_{F_{u}}\geq\tau_{G}$ $\check{H}_{0}:\tau_{F_{u}}\leq\tau_{G}$
$\epsilon=0.01$ $\epsilon=0.025$ $\alpha_{n}$ $\tilde{\alpha}_{n}$ $Q_{n}$ $T_{n}$
$\tau=240$ $\tau=360$ $\tau=240$ $\tau=360$
$\hat{f}_{nh}^{SG}$ $\hat{f}_{n}^{G}$ $\hat{f}_{nh}^{SG}$ $\hat{f}_{n}^{G}$ $\hat{f}_{nh}^{SG}$ $\hat{f}_{n}^{G}$ $\hat{f}_{nh}^{SG}$ $\hat{f}_{n}^{G}$
90 0.538 0.999 0.542 0.999 0.503 0.999 0.530 0.999 0.002 0.367 0.134 0.089
110 0.588 0.999 0.649 0.000 0.380 0.000 0.542 0.000 0.000 0.049 0.002 0.106
130 0.144 0.000 0.234 0.000 0.035 0.000 0.131 0.000 0.000 0.001 0.000 0.167
150 0.007 0.000 0.015 0.000 0.001 0.000 0.006 0.000 0.000 0.000 0.000 0.303
171 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
Table 1 reports the $p$-values of testing different hypotheses ($\tilde{H}_{0}$, $H_{0}$ or $\check{H}_{0}$) at different cutoffs. With a follow-up cutoff equal to the end of the study or 171 months, the proposed procedure (using $\hat{f}_{n}^{G}$ and $\hat{f}_{nh}^{SG}$) rejects $\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}$ at the 5% level when $\epsilon=0.01$ or 0.025, deciding the follow-up is sufficient for the two different $\tau$’s. The $\alpha_{n}$, $\tilde{\alpha}_{n}$, $Q_{n}$ tests also reject $H_{0}:\tau_{F_{u}}\geq\tau_{G}$ at the 5% level and hence in favor of sufficient follow-up. We note that the parameter $\epsilon^{\ast}$ for the $T_{n}$-test is set to $y_{(n)}$ since $2(y_{(n)}-\tilde{y}_{(n)})\geq y_{(n)}$ ($\epsilon^{\ast}$ instead of $\epsilon$ is used to avoid any confusion). If we consider $y_{(n)}$ and $\tilde{y}_{(n)}$ as the approximation of $\tau_{G}$ and $\tau_{F_{u}}$, respectively, then $\epsilon^{\ast}=y_{(n)}=171$ would possibly fall within the interval $(\tau_{G}-\tau_{F_{u}},2(\tau_{G}-\tau_{F_{u}})]\approx(91,182]$. Hence, the asymptotic distribution of the test statistic $T_{n}$ is degenerated as mentioned in Section 3 of Xie et al., (2023), while the $T_{n}$-test still concludes the follow-up as sufficient without applying the bootstrap procedure.
By shortening the follow-up period, all methods possess a monotonicity behavior of deciding sufficient follow-up. For example, the proposed procedure using $\hat{f}_{nh}^{SG}$ with $\tau=240$ and $\epsilon=0.01$ decides sufficient follow-up at the 5% level when the cutoff is at or after 150 months, while changes to conclude insufficient follow-up when the cutoff is less than 150 months. In terms of the cutoff at which the test decision changed, for example 150 for $\hat{f}_{nh}^{SG}$ and 130 for $\hat{f}_{n}^{G}$ when $\epsilon=0.01$, the procedure using $\hat{f}_{n}^{G}$ is less conservative than that using $\hat{f}_{nh}^{SG}$. This behavior was also observed in the simulation study. The procedure using $\hat{f}_{n}^{G}$ or $\hat{f}_{nh}^{SG}$ is more conservative than the $\alpha_{n}$, $\tilde{\alpha}_{n}$ and $Q_{n}$ tests for this dataset. It was also observed in the simulation study, e.g. in Setting 1, that $\alpha_{n}$, $\tilde{\alpha}_{n}$ and $Q_{n}$ tests exhibit higher rejection rate when $\tau_{G}$ is slightly greater than $q_{1-\epsilon}$. For the choice of $\tau$, a larger $\tau$ results in a more conservative test procedure, as shown in the simulation results. However, the conclusion does not change significantly when $\tau$ increases from 240 to 360. For the $T_{n}$-test, we can observe that it does not reject sufficient follow-up and the $p$-values decrease from 0.3 to around 0.09 as the follow-up cutoff reduces from 150 to 90.
5.2 Breast cancer study II
We consider a dataset of breast cancer patients extracted from the Surveillance, Epidemiology and End Results (SEER) database. The database ‘Incidence-SEER Research Data, 8 Registries, Nov 2022 Sub (1975-2020)’ with follow-up until December 2020 was selected and the breast cancer patients diagnosed in 1992 were extracted. This allows a maximum of 347 months (about 29 years) of follow-up. We excluded the observations with unknown or zero follow-up time, with unstaged or unknown cancer stage, and restricted the dataset to white patients with ages less than 60 and with Grade II tumor grade at diagnosis. The event time of interest is the time to death because of breast cancer. This cohort consists of 1233 observations with follow-up ranging from 1 to 347 months and has a censoring rate of 75.91%. A similar breast cancer dataset extracted from the SEER database was also studied by Tai et al., (2005), which suggested the minimum required follow-up time for Grade II breast cancer is 26.3 years if a log-normal distributed uncured survival time is assumed.
Figure 6: Kaplan-Meier estimate (solid) and its least concave majorant (dotted) for the SEER data.
Figure 6 depicts the Kaplan-Meier curve for the breast cancer data, which shows a short plateau. The maximum uncensored survival time $\tilde{y}_{(n)}$ and the maximum observed survival time $y_{(n)}$ are 341 and 347 months, respectively, and there are 295 censored observations on the plateau. Such a visual inspection based on the Kaplan-Meier curve is not adequate to assess sufficient follow-up. We apply the introduced testing procedure (using $\hat{f}_{n}^{G}$ and $\hat{f}_{nh}^{SG}$) and the four methods ($\alpha_{n}$, $\tilde{\alpha}_{n}$ and $Q_{n}$) discussed in Section 2 to this dataset for testing the null hypothesis of insufficient follow-up. We do not apply the $T_{n}$ test to this dataset given that a short plateau is observed from the Kaplan-Meier curve and there is no strong evidence of sufficient follow-up. Besides using the data with a follow-up cutoff in December 2020, we again construct also hypothetical scenarios with shorter follow-up. For the proposed procedure (using $\hat{f}_{n}^{G}$ and $\hat{f}_{nh}^{SG}$), the Kaplan-Meier curve in Figure 6 supports the assumption of decreasing density in the tail, at least starting from around 70 months. We again use two choices of $\tau$, namely 360 and 480 months. This means that we can consider the probability of a cancer related death after 30 or 40 years as negligible.
Table 2: $p$-values of testing $\tilde{H}_{0}$ or $H_{0}$ at different follow-up cutoffs for the SEER dataset (rounded to three decimal places). Bold indicates rejecting the null at the 5% level.
Cutoff $\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}$ $H_{0}:\tau_{F_{u}}\geq\tau_{G}$
$\epsilon=0.01$ $\epsilon=0.025$ $\alpha_{n}$ $\tilde{\alpha}_{n}$ $Q_{n}$
$\tau=360$ $\tau=480$ $\tau=360$ $\tau=480$
$\hat{f}_{nh}^{SG}$ $\hat{f}_{n}^{G}$ $\hat{f}_{nh}^{SG}$ $\hat{f}_{n}^{G}$ $\hat{f}_{nh}^{SG}$ $\hat{f}_{n}^{G}$ $\hat{f}_{nh}^{SG}$ $\hat{f}_{n}^{G}$
180 0.991 0.999 0.992 0.999 0.988 0.999 0.990 0.999 0.368 0.368 0.75
200 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.050 0.135 0.316
220 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.050 0.135 0.563
240 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.368 0.368 0.563
260 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.135 0.368 0.422
280 0.881 0.999 0.906 0.999 0.788 0.999 0.881 0.999 0.135 0.368 0.563
300 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.368 0.368 0.75
320 0.999 0.999 0.999 0.999 0.998 0.999 0.999 0.999 0.018 0.135 0.422
340 0.875 0.999 0.983 0.999 0.305 0.999 0.969 0.999 0.999 0.368 0.75
347 0.739 0.999 0.947 0.999 0.007 0.001 0.931 0.999 0.135 0.368 0.563
Table 2 reports the $p$-values of testing different hypotheses ($\tilde{H}_{0}$ or $H_{0}$) of sufficient follow-up at different follow-up cutoffs. At 5% level and with a follow-up cutoff in December 2020 (347 months), the proposed test procedure (using $\hat{f}_{n}^{G}$ and $\hat{f}_{nh}^{SG}$) does not reject $\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}$ when $\epsilon=0.01$, which decides that the follow-up is insufficient. However, the follow-up is considered sufficient when considering a more relaxed notion of sufficient follow-up for $\epsilon=0.025$ and $\tau=360$. The $\alpha_{n}$, $\tilde{\alpha}_{n}$ and $Q_{n}$ tests do not reject $H_{0}:\tau_{F_{u}}\geq\tau_{G}$ at the 5% level and conclude in favor of insufficient follow-up. If the follow-up period is shortened, all tests apart from $\alpha_{n}$ do not reject the hypothesis of insufficient follow-up. Moreover, for the proposed procedure the decision remains the same for all considered choices of $\epsilon$ and $\tau$.
6 Conclusion and discussion
In this paper, we proposed a new test for the null hypothesis of insufficient follow-up based on a more relaxed and realistic notion of ‘practically’ sufficient follow-up, meaning that the probability for the event to happen after the end of the study is smaller than a prespecified threshold. The test relies on the assumption that the density function of the survival times is non-increasing in the tail region two estimators under such shape constrained are considered. Despite guarantees on the asymptotic level of the test, we observe that in practice the test based on the smoothed Grenander estimator and a bootstrap procedure behaves the best. An extensive simulation study showed that the proposed procedure performs well in terms of both level and power, outperforming most of the time the existing methods. However, we warn that for scenarios with very high censoring rate, low uncured fraction and low sample size the empirical level is larger than the nominal one, but this is a problem even for the other existing tests. The choice of $\tau$ for the proposed method, such that it is almost impossible for the event to happen after $\tau$, should rely on medical knowledge. When such information is not available or uncertain, we suggest to use a larger $\tau$ for a more conservative test and to perform a sensitivity analysis on $\tau$. Furthermore, an interesting extension of the proposed test would be to consider testing sufficient follow-up in the presence of covariates.
Appendix A Appendix
Technical lemmas and proofs
Recall that $\hat{F}_{n}$ is the Kaplan-Meier estimator of the distribution function $F$. Assume that $F$ is continuous and G is right-continuous. Major and Rejtö, (1988) showed that the process $\hat{F}_{n}(t)-F(t)$ can be approximated by a Gaussian process on the time interval $[0,t_{0}]$ with $t_{0}<\tau_{H}$. Note that the right extreme of the observed event time $\tau_{H}$ can be expressed as $\tau_{H}=\tau_{F}\wedge\tau_{G}$. Note that $\tau_{F}=\infty$ in the presence of cured subjects. We therefore only consider the time up to $\tau_{G}$. With an additional assumption on the censoring distribution that $G$ has a jump at its right extreme $\tau_{G}$, such result can be extended to the time interval $[0,\tau_{G}]$. This can be shown using the construction in Remark 3 of Major and Rejtö, (1988).
Lemma 1.
Suppose the censoring distribution $G$ has a jump at $\tau_{G}$, i.e. $\Delta G(\tau_{G})=1-G(\tau_{G}-)>0$. The we have, for $x>0$,
$\mathbb{P}\left[\sup_{t\leq\tau_{G}}{n}\left|\hat{F}_{n}(t)-F(t)-n^{-1/2}\left\{1-F(t)\right\}W\circ L(t)\right|>K_{1}\log n+x\right]<K_{2}e^{-K_{3}x},$
where $K_{1}$, $K_{2}$ and $K_{3}$ are positive constants, $W$ is a Brownian motion, and
$L(t)=\int_{0}^{t}\frac{\mathrm{d}{F}(u)}{\left(1-G(u-)\right)\left(1-F(u)\right)^{2}}.$
With Lemma 1, we apply the arguments in Lopuhaä and Musta, (2017) to show the asymptotic normality of $\hat{f}_{nh}^{SG}(\tau_{G})$ below. We denote the least concave majorant of the Kaplan-Meier estimator by $\hat{F}_{n}^{G}$.
Proof of Theorem 2.
To show the asymptotic normality of $\hat{f}_{nh}^{SG}(\tau_{G})$, we first decompose $\hat{f}_{nh}^{SG}(\tau_{G})-f(\tau_{G})$ into three parts:
$\begin{split}\hat{f}_{nh}^{SG}(\tau_{G})-f(\tau_{G})&=\int_{\tau_{G}-h}^{\tau_{G}}\frac{1}{h}k_{B,\tau_{G}}\left(\frac{\tau_{G}-u}{h}\right)\mathrm{d}{F(u)}-f(\tau_{G})\\ &\quad+\int_{\tau_{G}-h}^{\tau_{G}}\frac{1}{h}k_{B,\tau_{G}}\left(\frac{\tau_{G}-u}{h}\right)\mathrm{d}{(\hat{F}_{n}-F)(u)}\\ &\quad+\int_{\tau_{G}-h}^{\tau_{G}}\frac{1}{h}k_{B,\tau_{G}}\left(\frac{\tau_{G}-u}{h}\right)\mathrm{d}{(\hat{F}_{n}^{G}-\hat{F}_{n})(u)}\\ &=(I)+(II)+(III).\end{split}$
For $(I)$, since $f_{u}$ is twice continuously differentiable and by the properties of the boundary kernel, we have
$\begin{split}n^{2/5}(I)&=\int_{\tau_{G}-h}^{\tau_{G}}\frac{1}{h}k_{B,\tau_{G}}\left(\frac{\tau_{G}-u}{h}\right)\left\{f(u)-f(\tau_{G})\right\}\mathrm{d}{u}\\ &=\int_{0}^{1}\frac{1}{h}k_{B,\tau_{G}}(v)\left\{f(\tau_{G}-hv)-f(\tau_{G})\right\}\mathrm{d}{v}\\ &=\int_{0}^{1}\frac{1}{h}k_{B,\tau_{G}}(v)\left\{-f^{\prime}hv+\frac{1}{2}f^{\prime\prime}(\xi_{n})h^{2}v^{2}\right\}\mathrm{d}{v}\\ &\xrightarrow{}\frac{1}{2}c^{2}f^{\prime\prime}(\tau_{G})\int_{0}^{1}v^{2}k_{B,\tau_{G}}(v)\mathrm{d}{v},\quad\text{as}~{}n\xrightarrow{}\infty,\end{split}$
where $00<\tau_{G}-\xi_{n}<hv<h\xrightarrow{}$ 0 as $n\xrightarrow{}0$.
For $(III)$, with the strong approximation in Lemma 1, we can apply the argument in the proof of Lemma 4.3 in Lopuhaä and Musta, (2017) to show that
$\sup_{t\in[0,\tau_{G}]}\left|\hat{F}_{n}^{G}(t)-\hat{F}_{n}(t)\right|=O_{P}\left(\frac{\log n}{n}\right)^{2/3}.$
The idea is to verify the four conditions in Durot and Lopuhaä, (2014) to establish such result. Therefore, $n^{2/5}(III)$ converges to zero in probability.
For $(II)$, using the strong approximation in Lemma 1, we can apply the argument in the proof of Theorem 4.4 in Lopuhaä and Musta, (2017) to show that $n^{2/5}(II)$, after a change of variable, is dominated by the term
$\frac{1}{\sqrt{hn^{1/5}}}\int_{0}^{1}k_{B,\tau_{G}}(v)\mathrm{d}{W_{n}(v)},$
where $W_{n}(v)=h^{-1/2}\left(1-F(\tau_{G})\right)\left\{W\circ L(\tau_{G}-hv)-W\circ L(\tau_{G})\right\}$, for $v\in[0,1]$. By scaling, time reversal, and symmetry of a Brownian motion, we have
[…]
where $\tilde{W}$ is a one-sided Brownian motion. By the uniform continuity of the one-sided Brownian motion on the compact interval $[0,1]$, we have
$\sup_{v\in[0,1]}\left|\tilde{W}\left(\frac{L(\tau_{G})-L(\tau_{G}-hv)}{h}\right)-\tilde{W}\left(L^{\prime}(\tau_{G})v\right)\right|\xrightarrow{\mathbb{P}}0,$
where $L^{\prime}(\tau_{G})=\frac{f(\tau_{G})}{\left(1-G(\tau_{G}-)\right)\left(1-F(\tau_{G})\right)^{2}}$. Therefore we have
$\left(1-F(\tau_{G})\right)\sqrt{L^{\prime}(\tau_{G})}\int_{0}^{1}k_{B,\tau_{G}}(v)\mathrm{d}{\tilde{W}(v)}\sim N\left(0,\frac{f(\tau_{G})}{1-G(\tau_{G}-)}\int_{0}^{1}k_{B,\tau_{G}}(v)^{2}\mathrm{d}{v}\right).$
Hence
[…]
Proof of Proposition 1.
Let $\tau_{G}\leq q_{1-\epsilon}$. We have
$\begin{split}&\mathbb{P}\left[\hat{f}_{n}^{G}(\tau_{G}-cn^{-a})\leq\frac{\epsilon\hat{F}_{n}(\tau_{G})}{\tau-\tau_{G}}-A_{1}^{-1}n^{-(1-a)/2}Q_{1-\alpha}^{G}\right]\\ &=\mathbb{P}\left[f(\tau_{G})-\hat{f}_{n}^{G}(\tau_{G}-cn^{-a})\geq f(\tau_{G})-\frac{\epsilon\{\hat{F}_{n}(\tau_{G})-F(\tau_{G})\}}{\tau-\tau_{G}}-\frac{\epsilon F(\tau_{G})}{\tau-\tau_{G}}\right.\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad+A_{1}^{-1}n^{-(1-a)/2}Q_{1-\alpha}^{G}\Bigg{]}\\ &\leq\mathbb{P}\left[A_{1}n^{(1-a)/2}\{f(\tau_{G})-\hat{f}_{n}^{G}(\tau_{G}-cn^{-a})\}\right.\\ &\qquad\qquad\geq Q_{1-\alpha}^{G}+\frac{A_{1}n^{(1-a)/2}p\eta}{\tau-\tau_{G}}-\frac{\epsilon A_{1}n^{(1-a)/2}\{\hat{F}_{n}(\tau_{G})-F(\tau_{G})\}}{\tau-\tau_{G}}\Bigg{]}.\end{split}$
Here the last inequality follows from the fact that $\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}$ entails $f(\tau_{G})\geq\frac{p(\epsilon-\eta)}{\tau-\tau_{G}}$, and $F(\tau_{G})\leq p$. By Theorem 1 and Lemma 1, the probability on the right hand side converges to $\alpha$, which concludes the proof. ∎
Proof of Proposition 2.
Let $\tau_{G}\leq q_{1-\epsilon}$. We have
$\begin{split}&\mathbb{P}\left(\hat{f}_{nh}^{SG}(\tau_{G})\leq\frac{\epsilon\hat{F}_{n}(\tau_{G})}{\tau-\tau_{G}}+n^{-2/5}Q_{\alpha}^{SG}\right)\\ &=\mathbb{P}\left(\hat{f}_{nh}^{SG}(\tau_{G})-f(\tau_{G})\leq\frac{\epsilon\{\hat{F}_{n}(\tau_{G})-F(\tau_{G})\}}{\tau-\tau_{G}}+\frac{\epsilon F(\tau_{G})}{\tau-\tau_{G}}-f(\tau_{G})+n^{-2/5}Q_{\alpha}^{SG}\right)\\ &\leq\mathbb{P}\left(\hat{f}_{nh}^{SG}(\tau_{G})-f(\tau_{G})\leq\frac{\epsilon\{\hat{F}_{n}(\tau_{G})-F(\tau_{G})\}}{\tau-\tau_{G}}+\frac{p\eta}{\tau-\tau_{G}}+n^{-2/5}Q_{\alpha}^{SG}\right)\\ &=\mathbb{P}\left(n^{2/5}\{\hat{f}_{nh}^{SG}(\tau_{G})-f(\tau_{G})\}\leq\frac{\epsilon n^{2/5}\{\hat{F}_{n}(\tau_{G})-F(\tau_{G})\}}{\tau-\tau_{G}}+\frac{n^{2/5}p\eta}{\tau-\tau_{G}}+Q_{\alpha}^{SG}\right).\end{split}$
Here the inequality follows from the fact that $\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}$ entails $f(\tau_{G})\geq\frac{p(\epsilon-\eta)}{\tau-\tau_{G}}$, and $F(\tau_{G})\leq p$. By Theorem 2 and Lemma 1, the probability on the right hand side converges to $\alpha$, which concludes the proof. ∎
Supplementary Material:
Testing for sufficient follow-up in survival data with a cure fraction
Tsz Pang Yuen and Eni Musta
Appendix S1 Simulation study
S1.1 Simulation settings
Table S1: Censoring rates for Setting 1.
$p$ $\Delta G(\tau_{G})$ Censoring rate
0.2 0 82.91% – 87.84%
0.02 82.85% – 87.73%
0.05 82.76% – 87.56%
0.2 82.34% – 86.69%
0.6 0 48.74% – 63.52%
0.02 48.56% – 63.16%
0.05 48.29% – 62.64%
0.2 47.02% – 60.04%
0.8 0 31.64% – 51.35%
0.02 31.39% – 50.87%
0.05 31.05% – 50.18%
0.2 29.33% – 46.72%
Table S2: Censoring rates for Setting 2.
$p$ $\lambda$ Censoring rate
0.2 0.4 91.13% – 91.18%
1 86.69% – 87.12%
5 81.86% – 83.30%
0.6 0.4 73.33% – 73.49%
1 60.02% – 61.29%
5 45.56% – 49.87%
0.8 0.4 64.46% – 64.66%
1 46.65% – 48.36%
5 27.34% – 33.09%
Table S3: Censoring rates for Setting 3.
$p$ $\Delta G(\tau_{G})$ Censoring rate
0.2 0 80.83% – 85.08%
0.02 80.82% – 85.02%
0.05 80.80% – 84.92%
0.2 80.68% – 84.46%
0.6 0 42.57% – 55.26%
0.02 42.53% – 55.08%
0.05 42.45% – 54.80%
0.2 42.08% – 53.41%
0.8 0 23.39% – 40.28%
0.02 23.32% – 40.03%
0.05 23.22% – 39.66%
0.2 22.73% – 37.81%
Table S4: Censoring rates for Setting 4.
$p$ $\Delta G(\tau_{G})$ Censoring rate
0.6 0 48.38% – 63.85%
0.02 48.22% – 63.48%
0.05 47.96% – 62.94%
0.2 46.71% – 60.28%
Table S5: Censoring rates for Setting 5.
$p$ $\lambda_{C}$ Censoring rate
0.6 0.5 45.24% – 49.75%
3 62.14% – 63.23%
S1.2 Simulation results
$q_{1}$, $q_{2}$, $q_{3}$, $q_{4}$, $q_{6}$, and $q_{12}$ are the 90%, 92.5%, 95%, 97.5%, 99%, and 99.9% quantiles of $F_{u}$, resp.; $q_{5}$ is the mid-point between $q_{4}$ and $q_{6}$; $q_{7},\cdots,q_{11}$ are 5 evenly separated points between $q_{6}$ and $q_{12}$ in the figures in this section. Table S6 reports their actual values for Settings 1–5.
Table S6: Actual values of $q_{1}$, …, $q_{12}$ for Settings 1–5.
$\tau_{G}$ Setting
1 2 3 4 5
$q_{1}$ 2.3026 0.4605 7.9528 2.2164 0.4433
$q_{2}$ 2.5903 0.5181 10.0642 2.4740 0.4948
$q_{3}$ 2.9957 0.5991 13.4616 2.8218 0.5644
$q_{4}$ 3.6889 0.7378 20.4117 3.3596 0.6719
$q_{5}$ 4.1470 0.8294 26.1116 3.6383 0.7277
$q_{6}$ 4.6052 0.9210 31.8114 3.9170 0.7834
$q_{7}$ 4.9889 0.9978 38.4388 4.0160 0.8032
$q_{8}$ 5.3727 1.0745 45.0661 4.1149 0.8230
$q_{9}$ 5.7565 1.1513 51.6935 4.2139 0.8428
$q_{10}$ 6.1402 1.2280 58.3209 4.3129 0.8626
$q_{11}$ 6.5240 1.3048 64.9483 4.4118 0.8824
$q_{12}$ 6.9078 1.3816 71.5756 4.5108 0.9022
S1.2.1 Setting 1
Figures S1S3 show the rejection rate against of insufficient follow-up $\tau_{G}$ for Setting 1, each figure with different $p$. To study the effect of the choice of $\tau$ on the rejection rate, we considered 3 different $\tau$’s, namely $\tau_{1}\approx 7.601$ (the 99.95% quantile of $F_{u}$), $\tau_{2}\approx 9.210$ (the 99.99% quantile of $F_{u}$), and $\tau_{3}=(\tau_{1}+\tau_{2})/2\approx 8.406$. The rejection rate of $\hat{f}_{nh}^{SG}$ using $\tau_{1}$, $\tau_{2}$, and $\tau_{3}$ are rendered by solid, long dashed, and dashed lines, respectively, in Figure S4. In summary, larger $\tau$ results in a more conservative procedure, i.e. better control on the empirical level but with some loss of empirical power.
Figure S1: Rejection rate of the null hypothesis of insufficient follow-up for different methods in Setting 1 when $p=0.2$ (uncured fraction).
Figure S2: Rejection rate of the null hypothesis of insufficient follow-up for different methods in Setting 1 when $p=0.6$ (uncured fraction).
Figure S3: Rejection rate of the null hypothesis of insufficient follow-up for different methods in Setting 1 when $p=0.8$ (uncured fraction).
Sensitivity of $\tau$
Figure S4: Rejection rate of the null hypothesis of insufficient follow-up for the test based on $\hat{f}_{nh}^{SG}$ with different $\tau$ in Setting 1 when $p=0.6$ (uncured fraction) .
S1.2.2 Setting 2
Figures S5S7 depict the rejection rate of insufficient follow-up against $\tau_{G}$ for Setting 2, each figure with different $p$. In Setting 2, the uncured subjects have the exponential distribution with rate $\lambda\in\{0.4,1,5\}$, while the censoring time has the exponential distribution with the rate fixed at 0.5. The censoring rate decreases as $\lambda$ increases, which affects the performance of the testing procedure. In particular, $\lambda=0.4$ corresponds to a very high censoring rate since the density of the censoring variable decreases quicker than that of the event times and all methods have problems in controlling the level of the test. We observe that each method performs better, in terms of empirical level and power, when $\lambda$ is larger.
Figure S5: Rejection rate for Setting 2 when $p=0.2$ (uncured fraction).
Figure S6: Rejection rate for Setting 2 when $p=0.6$ (uncured fraction).
Figure S7: Rejection rate for Setting 2 when $p=0.8$ (uncured fraction).
S1.2.3 Setting 3
Figures S8S10 depict the rejection rate of insufficient follow-up against $\tau_{G}$ for Setting 3 when $p$ is 0.2, 0.6, and 0.8. This setting the density $f_{u}$ decreases faster and we observe a worse behaviour in terms of empirical level, particularly when $p$, $n$ and $\Delta G(\tau_{G})$ are smaller.
Figure S8: Rejection rate for Setting 3 when $p=0.2$ (uncured fraction).
Figure S9: Rejection rate for Setting 3 when $p=0.6$ (uncured fraction).
Figure S10: Rejection rate for Setting 3 when $p=0.8$ (uncured fraction).
Changing $\epsilon$ from 0.01 to 0.005
Figure S11 shows the rejection rates of insufficient follow-up for $\hat{f}_{nh}^{SG}$, when $\epsilon=0.005$ is used for $\tilde{H}_{0}:\tau_{G}\leq q_{1-\epsilon}$, meaning that we consider the follow-up as insufficient when $\tau_{G}$ is less than or equal to the 99.5% quantile of $F_{u}$. Under such a stricter characterization of insufficient follow-up, the rejection proportion is lower compared to the results for $\epsilon=0.01$. Note that the parameter $\tau$ is fixed to the 99.95% quantile of $F_{u}$ for both cases.
Figure S11: Rejection rate of insufficient follow-up for the test based on $\hat{f}_{nh}^{SG}$ in Setting 3 when $p=0.6$ (uncured fraction). $q_{6}$ corresponds to $q_{0.99}$; and $q_{0.995}$ locates between $q_{7}$ and $q_{8}$.
S1.2.4 Setting 4
Figure S12: Rejection rate of insufficient follow-up for different methods in Setting 4 when $p=0.6$ (uncured fraction).
S1.2.5 Setting 5
Figure S13 shows the rejection rate of insufficient follow-up against $\tau_{G}$ for Setting 5. The uncured subjects have a truncated exponential distribution with parameter $5$ and the censoring time follows an exponential distribution with rate $\lambda_{C}\in\{0.5,3\}$ and truncated at $\tau_{G}$. The censoring becomes heavier when $\lambda_{C}$ increases. $\hat{f}_{nh}^{SG}$ has a steeper rejection curve among the investigated methods, although the rejection rate of $\hat{f}_{nh}^{SG}$ is higher than the significance level when $\tau_{G}$ is between $q_{4}$ and $q_{6}$, and $\lambda_{C}=3$. The $Q_{n}$ test shows very little power to detect sufficient follow-up, particularly when the censoring rate is higher.
Figure S13: Rejection rate Setting 5 when $p=0.6$ (uncured fraction).
S1.2.6 Setting 6
In Setting 6, the uncured subjects have the exponential distribution with rate of 1 and the censoring time follows the exponential distribution with rate of 0.5. In particular $\tau_{F_{u}}=\infty$ and $\tau_{G}=\infty$, and this setting is regarded as sufficient follow-up under $\check{H}_{0}:\tau_{F_{u}}\leq\tau_{G}$ but is it not sufficient follow-up under our formulation $\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}$. Table S7 shows the rejection proportion of null hypotheses considered by the different tests. For the $T_{n}$-test, for which $\check{H}_{0}$ is true, the empirical level is around the significance level. For computation of our test statistics we use the maximum observed survival time $y_{(n)}$ instead of $\tau_{G}$ for each generated dataset, meaning that essentially, we consider the follow-up as insufficient when $q_{1-\epsilon}\geq y_{(n)}$. For the simulation with $p=0.8$ and sample size of 200, there are 5 out of 500 replications that $q_{1-\epsilon}<y_{(n)}$. This is not observed for the remaining cases. The parameter $\tau$ is set to $y_{(n)}+\delta$, where $\delta\in\{1,2\}$, which is considered reasonable given the range of values of $y_{(n)}$. The empirical power of the proposed methods using $\hat{f}_{nh}^{SG}$ and $\hat{f}_{n}^{G}$ are close to 1, among all different $p$’s and sample sizes.
Table S7: Simulation result (in terms of rejection rate of the null hypotheses) for Setting 6. $\tau$ is set to $y_{(n)}+\delta$ for the tests based on $\hat{f}_{n}^{G}$ and $\hat{f}_{nh}^{SG}$.
$\tilde{H}_{0}:q_{1-\epsilon}\geq\tau_{G}$ ${H}_{0}:\tau_{F_{u}}\geq\tau_{G}$ $\check{H}_{0}:\tau_{F_{u}}\leq\tau_{G}$
$p$ size $\delta$ $\hat{f}_{n}^{G}$ $\hat{f}_{nh}^{SG}$ $\alpha_{n}$ $\tilde{\alpha}_{n}$ $Q_{n}$ $T_{n}$
0.2 200 1 1 1 0.998 0.814 0.996 0
2 1 1
500 1 1 1 0.994 0.854 0.994 0.002
2 1 1
1000 1 1 1 1 0.898 1 0.002
2 1 1
0.6 200 1 0.992 0.988 0.976 0.734 0.956 0
2 0.992 0.984
500 1 1 1 0.992 0.79 0.98 0
2 1 0.998
1000 1 1 1 0.994 0.876 0.99 0
2 1 1
0.8 200 1 0.974 0.942 0.948 0.624 0.876 0.008
2 0.974 0.928
500 1 0.996 0.994 0.97 0.728 0.94 0.014
2 0.996 0.986
1000 1 0.998 0.998 0.98 0.81 0.966 0.008
2 0.998 0.998
Appendix S2 Algorithm
Algorithm S1 Bootstrapping procedure
1:
2:
Orignal sample data
$\left\{\left(y_{i},\delta_{i}\right):i=1,\cdots,n\right\}$
Level of the test
$\alpha$
Number of bootstrap iterations
$B$
3:Obtain a smooth KME of $F$ by
$\tilde{F}_{nh_{0}}(t)=\int_{(t-h_{0})\vee 0}^{(t+h_{0})\wedge\tau_{G}}\frac{1}{h_{0}}k^{(t)}\left(\frac{t-v}{h_{0}}\right)\hat{F}_{n}^{G}(v)\mathrm{d}{v},\quad t\in[0,\tau_{G}],$
where $\hat{F}_{n}^{G}$ is the least concave majorant (LCM) of the KME $\hat{F}_{n}$.
4:Compute the derivative $\tilde{f}_{nh_{0}}(\tau_{G})$ of $\tilde{F}_{nh_{0}}$ at $\tau_{G}$.
5:Estimate $G$ by the reversed KME, denoted by $\hat{G}_{n}$, using $\left\{(y_{i},\delta_{i}):i=1,\cdots,n\right\}$.
6:for $b=1,\cdots,B$ do
7:     Draw $(c_{b,1}^{\ast},\cdots,c_{b,n}^{\ast})$ from $\hat{G}_{n}$;
8:     Draw $(t_{b,1}^{\ast},\cdots,t_{b,n}^{\ast})$ from $\tilde{F}_{nh_{0}}$;
9:      $\triangleright$ Set $t_{b,i}^{\ast}=\infty$ if $u_{b,i}>\sup_{t}\tilde{F}_{nh_{0}}(t)$, where $u_{b,i}$ is the generated standard uniform random variate.
10:     Construct $\{{({y_{b,i}^{\ast},\delta_{b,i}^{\ast}}),i=1,\cdots,n}\}$, where $y_{b,i}^{\ast}=t_{b,i}^{\ast}\wedge c_{b,i}^{\ast}$ and $\delta_{b,i}^{\ast}=\mathds{1}_{\{{t_{b,i}^{\ast}\leq c_{b,i}^{\ast}}\}}$.
11:     Obtain the bootstrap estimate $\hat{f}_{nh,b}^{SG^{\ast}}(\tau_{G})$ using $\{{({y_{b,i}^{\ast},\delta_{b,i}^{\ast}}):i=1,\cdots,n}\}$.
12:end for
13:Approximate the critical value of the test by the ${\alpha}$-quantile of $\{\hat{f}_{nh,b}^{SG^{\ast}}(\tau_{G})-\tilde{f}_{nh_{0}}(\tau_{G}):b=1,\cdots,B\}$